OpenStudy (anonymous):

you solve the equation x/9 = -7 by multiplying both sides by 9. You conclude that the value of x is -63. Explain whether this situation represents inductive reasoning or deductive reasoning

OpenStudy (anonymous):

@Lyssa123

OpenStudy (anonymous):

it would look like this 9* -63/9=-7 *9

OpenStudy (anonymous):

so its deductive reasoning ? @Lyssa123

OpenStudy (anonymous):

I can't u that yet until we do the math

OpenStudy (anonymous):

ok start

OpenStudy (anonymous):

ok hold on a min

OpenStudy (anonymous):

so its -63 = 63 @Lyssa123

OpenStudy (anonymous):

no , it say to multiply 9 on both sides 9* -63 = -567 and -7*9=-63

OpenStudy (anonymous):

so inductive or deductive? @Lyssa123

OpenStudy (anonymous):

deductive is the answer